DX AO Observations

08:46 Starting Authorization
  • Shell set with only 140 modes
08:58
  • Preset Sent
  • Stated Mag 14.84
  • Binning 4
  • 36 Modes
  • 200 Hz
09:07
  • Found Proper Guide start
  • Estimated magnitude 15.3
  • Seeing 1.0
  • Gains 3.40 3.00 3.00
  • Optical Loop Data 2000 frames
  • 20170414_091242
09:16
  • Barry is having scheduler issues.
  • It turned out to be a problem with GEIRS
09:30
  • Resuming
  • Optical loop data
  • 20170414_093033
09:35
  • Loop Opened
09:52
  • Resending preset on both sides
  • Bin 4
  • 200 Hz
  • 36 Modes
  • Gains: 3.40 1.10 1.10
  • Optical Loop Data 2000 Frames
  • 20170414_100024
  • 20170414_100646
  • 20170414_101720
  • 20170414_103744
11:01
  • SX Failure
11:17
  • Preparing to change to new target.
  • Doing Pointing Check
11:25
  • Preseting to new object
  • Stated Magnitude 13.8
  • Starting in bin 3
  • Changing to bin 4
  • 36 modes
  • 200 Hz
  • Gains: 3.40 3.00 3.00
  • Dataset
  • 20170414_113156
  • 20170414_113247
  • Dimm 1.30
  • Big vibrations, AO winds high. TTOVMS 51mas X rms, 75mas Y rms
  • 20170414_114601
12:09
  • Preset to PSF star
  • Bin 3
  • 64 modes
  • 312 Hz
  • Listed mag 13.35
  • Mag 13.1 BC of bacground
  • Gains:
  • This offset failed, it is now too bright.
-- %USERSIG{GregoryTaylor - 2017-04-14}%
Topic revision: r3 - 14 Apr 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback